Healthtrax8:

How do you figure out the percentage without a calculator? In store stock to 152 cans of popcorn for the weekend sale they sold 72 what percentage was sold?

Gdeinward:

do you know how to calculate percentage?

Healthtrax8:

By 0.25

Shadow:

You could do long division. 6/10 = 0.6 -> 60% 6 is 60% of 10. Therefore doing 72/152 should get you a decimal that you can convert to the percentage that 72 is of 152. Does that make sense?

Healthtrax8:

Yes

Healthtrax8:

I’m taking an exam for placement today haven’t study math in 15 years

Shadow:

I see. Placement in what, if you don't mind me asking?

Healthtrax8:

60% of what number is equal to 30? How would I figure that problem out

Healthtrax8:

My answer was 18

Vocaloid:

60% as a decimal is 60/100 = 0.6 “Of” means multiply “A number” can be represented as some variable x So translating the problem into an equation 0.60 * x = 30 You can divide both sides by 0.60 to get x